OpenStudy (anonymous):

How to find the exact value of 6pi without a calculator

myininaya (myininaya):

you can't 6pi is irrational it would only give you up to a certain amount of digits

myininaya (myininaya):

what i'm saying 6pi is the exact value the calculator will give you an approximation

OpenStudy (anonymous):

But, you can get arbitrary close to the exact value by hand, using this formula: \[\frac{\pi}{4} = 1-\frac{1}{3}+\frac{1}{5}-\ldots\]